The residents of the town Ectenia all love economics, and the mayor proposes building an economics museum. The museum has a fixed cost of $2,400,000 and no variable costs. There are 100,000 town residents, and each has the same demand for museum visits: 𝑄𝐷=10−𝑃 where P is the price of admission.

1
Expert's answer
2021-12-21T12:08:39-0500

Solution:

Given:

Fixed cost = FC = $2,400,000

Number of residents = N = 100,000

Average Fixed Cost = AFC = "\frac{FC}{N}""\frac{2,400,000}{100,000}=" $24

FCN = $24


Since there is no variable cost,

Marginal Cost = MC = 0





Figure 1.

It is a monopoly market as the museum is constructed and regulated by the government body only. No private sector has any power over its construction and regulation.

b. As the mayor charges no price for the entry, the price is zero. The equation gives the demand curve for the firm,

QD=10-P,

So, as the price of the museum visit is zero, the number of visits for each person is 10 visits.

As it is a monopoly market, the price equals the average cost, which is $24, then the consumer surplus is,

=$24×10=$240

The benefit each person would get

=Consumer surplus−tax= $240−$24 = $216

=$216

Thus, each person gets a benefit of $216.